# Bayesian neural networks with interpretable priors from Mercer kernels

### v1

@@ Abstract
Quantifying the uncertainty in the output of a neural network is essential for deployment in scientific or engineering applications where decisions must be made under limited or noisy data. Bayesian neural networks (BNNs) provide a framework for this purpose by constructing a Bayesian posterior distribution over the network parameters. However, the prior, which is of key importance in any Bayesian setting, is rarely meaningful for BNNs. This is because the complexity of the input-to-output map of a BNN makes it difficult to understand how certain distributions enforce any interpretable constraint on the output space. Gaussian processes (GPs), on the other hand, are often preferred in uncertainty quantification tasks due to their interpretability. The drawback is that GPs are limited to small datasets without advanced techniques, which often rely on the covariance kernel having a specific structure. To address these challenges, we introduce a new class of priors for BNNs, called Mercer priors, such that the resulting BNN has samples which approximate that of a specified GP. The method works by defining a prior directly over the network parameters from the Mercer representation of the covariance kernel, and does not rely on the network having a specific structure. In doing so, we can exploit the scalability of BNNs in a meaningful Bayesian way.
Keywords: Bayesian neural networks, Gaussian processes, reproducing kernel Hilbert spaces, inverse problems
@@ 1 Introduction
As neural networks begin to be employed in sensitive areas, such as complex engineering systems or healthcare [55, 44, 27, 21], the need to reliably assess their predictions becomes critically important. In these settings, it is oftentimes insufficient for a neural network to provide accurate point predictions, as the model must also quantify the uncertainty in the output to inform decisions. For neural networks, this task can be formulated under the Bayesian paradigm by treating the network parameters as random variables and constructing a posterior distribution conditioned on the data, which leads to a Bayesian neural network (BNN) [31, 36]. The complexity of the input-to-output map of a neural network makes it difficult to enforce any interpretable constraints on the output space of the BNN when assigning a specific form of the prior. As a result, the default choice is simply to place an independent and identically distributed Gaussian prior over the individual network parameters. Although this choice is certainty convenient, it is not always appropriate, and rarely translates to any meaningful constraints. In contrast, Gaussian processes (GPs), offer a greater degree of interpretability for uncertainty quantification tasks. This comes at a greater computational cost, however, as GPs do not scale well to large datasets without clever implementations that restrict the possible covariance structures.
Notably, a close relationship between BNNs and GPs is well-documented in the literature [29, 34, 38, 14]. For example, when the parameters of a BNN are drawn independently from a Gaussian distribution in the usual manner, the resulting distribution over functions converges to a GP in the infinite-width limit [36]. However, the form of the limiting GP depends entirely from the activation function of the network, as this determines the covariance kernel. This means that specifying a desired GP prior would require identifying a tailor-made activation function whose induced kernel matches that of the GP. While this has been shown to be possible for certain cases [58, 29, 64, 43], a general procedure remains out of reach. This observation motivates an alternative viewpoint: rather than fixing the activation function and sampling the parameters independently, is it instead possible to design the parameter distribution so that the BNN approximates a given GP?
@@
We first motivate the technique and contextualize the Mercer prior within the available methods for sampling BNNs in Sec. 2. The prior is introduced in Sec. 3, where we also derive a scalable sampling scheme and discuss possible covariance choices. After defining the Mercer prior, we do a deep dive into its use in sampling BNNs which resemble Brownian motion draws in Sec. 4. We compare the output statistics of the BNN to true Brownian motion, and provide numerical evidence that convergence is expected in the infinite-width limit. We then present three application-driven examples in Sec. 5: (i) hierarchical GP regression under heteroscedastic noise, (ii) time series prediction on data with a periodic structure, and (iii) an elliptic PDE inverse problem with real-world data. Together, these results illustrate how Mercer priors provide a principled and scalable way to endow BNNs with an interpretable structure, enabling them to be deployed in uncertainty quantification tasks that are currently out of reach for standard GP models.
@@ 2 Background and setting
$u_{\theta}(x)=\sum_{j=1}^{N_{L}}w^{L}_{1,j}h\left(z_{j}^{L}(x)\right),\quad z_{i}^{s}(x)=\sum_{j=1}^{N_{s-1}}w_{i,j}^{s-1}h\left(z_{j}^{s-1}(x)\right)+b_{i}^{s-1},\quad s=2,\dots,L,$ (2.1)
where the input layer follows $z_{i}^{1}(x)=\sum_{j=1}^{d}w_{i,j}^{0}x_{j}+b_{i}^{0}$. We let $N_{s}=\dim(z^{s}(x))$ denote the number of neurons in layer $s$. The parameters of the BNN are the weights $w_{i,j}^{s}\in\mathbb{R}$ and the biases $b_{i}^{s}\in\mathbb{R}$, which are consolidated into the collection $\theta$. Eq. (2.1) defines a stochastic process $u:\Omega\times\Theta\to\mathbb{R}$, where $\Theta$ is a measurable parameter space on which a prior distribution on the BNN parameters $\theta$ is assigned. In the case $\dim(\Theta)<\infty$, i.e., for a finite-width network, we denote the prior over the BNN parameters as $p(\theta)$. We are also interested in studying infinite-width networks, where it is no longer rigorous to write $p(\theta)$. In this case, we denote the prior as some probability measure $\mu_{0}$ over the space $\Theta$. Either way, the mapping $\theta\mapsto u_{\theta}$ is a random field taking values in $L^{2}(\Omega)$.
@@
which holds even when $\dim(\Theta)=\infty$ [50]. In the case of a finite-width network, this reduces to the well-known formula $p(\theta|y)\propto p(y|\theta)p(\theta)$, with the likelihood being $p(y|\theta)\propto\exp\left(-\tfrac{1}{2}\|\Gamma^{-1}(y-R(u_{\theta}))\|^{2}\right)$.
@@ 2.1 Wide neural networks and Gaussian processes
$k(x,x^{\prime})=\sigma_{w}^{2}\mathbb{E}[h(x^{T}w_{1}^{0}+b_{1}^{0})h(x^{\prime T}w_{1}^{0}+b_{1}^{0})]+\sigma_{b}^{2}.$
@@
$u(x)=\sum_{n=1}^{\infty}\sqrt{\lambda_{n}}\xi_{n}\phi_{n}(x),$
where $\lambda_{n}$ and $\phi_{n}$ are the eigenvalues and eigenfunctions of the kernel integral operator $Su=\int k(\cdot,x^{\prime})u(x^{\prime})dx^{\prime}$, and $\xi_{n}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,1)$. The expansion is truncated to a finite number of terms $M$ in practice. When using the KLE, the computational complexity is transferred from inverting a large matrix to constructing the basis functions. However, since the eigenfunctions $\phi_{n}$ are global and fixed, performance may suffer when the data is non-uniformly spaced or sparse, unless $M$ is large. Also, the KLE suffers from the curse of dimensionality, as the number of eigenfunctions needed to capture most of the variance grows exponentially with the dimension. A comprehensive review of these ideas can be found in [30].
@@ 2.2 Related methods
Looking to gain both the expressivity of GPs and scalability of BNNs, several methods have been proposed that exploit the relationship between the two. This can be done using the correspondence between a BNN and the associated infinite-width limit GP, so that the problem of prior selection for a BNN is transferred to the easier problem of prior selection for GPs. One class of methods selects the prior in order to minimize some objective between the BNN and the target GP [18, 20, 51]. The drawback is that the user is constrained to a particular algorithm for posterior inference (like variational inference). These methods also typically need to see some of the training data upfront in order to construct the prior, e.g. [35]. Other methods rely on changing the architecture of the BNN to ensure convergence to the GP in an appropriate sense [41, 2, 43]. This of course works on a case-by-case basis.
@@
The ridgelet prior and the Mercer prior we present in this work are attempting to solve the same problem. Namely that we want to sample the parameters of a BNN such that the resulting function resembles draws from a predetermined GP. The difference then comes down to how the prior over the network parameters is defined. In this sense, one can view the two as different finite-dimensional approximations of the same GP. However, computing the Mercer prior scales linearly in the number of network parameters, which allows for the potential use of larger networks. While the Mercer prior does contain two integrals which need to be computed, we avoid the curse of dimensionality by deriving a sampling scheme which uses an unbiased estimate of the prior. This is built with Monte Carlo estimates, which in principle guarantees convergence even for a batch size of $1$ (this is equivalent to setting $D=1$ in the ridgelet prior).
@@ 3 The Mercer prior
Start by defining a GP on $\Omega\subset\mathbb{R}^{d}$ that we wish to emulate $u\sim\mathcal{GP}(0,k)$, where $k:\Omega\times\Omega\to\mathbb{R}$ is the covariance kernel. Note that we are choosing to work with a centered GP to simplify some of the details. A non-zero mean function could be used instead, provided that it satisfies certain regularity constraints. The resulting formulae would then need to be adjusted. Alternatively, one could shift the space so that the GP becomes centered. Further, for the covariance kernel, we enforce the restriction $\int_{\Omega}k(x,x)dx<\infty$. This is so that we may use the following theorem, which relates the GP to an induced Gaussian measure on $L^{2}(\Omega)$. A positive-definite kernel statisfying these properties is sometimes called a Mercer kernel.
@@
$\int_{\Omega}m^{2}(x)dx<\infty,\quad\int_{\Omega}k(x,x)dx<\infty.$
In this case, $u$ induces the Gaussian measure $\mathcal{N}(m,S)$ on $L^{2}(\Omega)$ with the covariance operator being $(Sv)(\cdot)=\int_{\Omega}k(\cdot,x)v(x)dx$, for $v\in L^{2}(\Omega)$.
@@
We then numerically characterize eq. (3.2) using traditional Monte Carlo sampling schemes in order sample BNNs. Each sample of $\theta$ provides a new neural network, and we will see that the ensemble statistics approximately follow that of the original Gaussian measure $\mu$. This offers a significant reduction in computational complexity when the BNN sampled from eq. (3.2) replaces a GP for the same task. As mentioned previously, the complexity of a GP scales with the cube of the domain mesh. On the other hand, the BNN prior of eq. (3.2) grows with $\theta$. After sampling, the BNN can be evaluated on a domain with a near-arbitrarily fine mesh, which allows for easy super-resolution.
@@
$(S^{-1}v)(\cdot)=\int_{\Omega}k^{-1}(\cdot,x)v(x)dt,\quad v\in L^{2}(\Omega),$
such that $S^{-1}(Su)=u$.
We work our way around the need to evaluate the integrals by deriving a sampling scheme which uses stochastic gradient Langevin dynamics (SGLD) [57]. The beauty of SGLD is that samples from $p(\theta)$ can be generated using only unbiased estimates of $\log\nabla_{\theta}p(\theta)=-1/2\nabla_{\theta}\langle u_{\theta},S^{-1}u_{\theta}\rangle$. This allows us to approximate the various integrals required using sampling averages and SGLD guarantees convergence to the correct probability distribution [53].
@@
$(Su)(\cdot)=\sum_{n=1}^{\infty}\lambda_{n}\langle u,\phi_{n}\rangle\phi_{n},\quad u\in L^{2}(\Omega),$
@@
Let $k:\Omega\times\Omega\to\mathbb{R}$ be a continuous, positive-definite kernel, and $S$ be the operator on $L^{2}(\Omega)$ defined by $u\mapsto\int_{\Omega}k(\cdot,x)u(x)dx.$ Denote the eigenvalues and eigenfunctions of $S$ as $\{\lambda_{n}\}_{n\in\mathbb{N}}$ and $\{\phi_{n}\}_{n\in\mathbb{N}}$, respectively. Then $k$ can be expressed as
$k(x,x^{\prime})=\sum_{n=1}^{\infty}\lambda_{n}\phi_{n}(x)\phi_{n}(x^{\prime}),\quad x,x^{\prime}\in\Omega,$
@@
$k^{-1}(x,x^{\prime})=\sum_{n=1}^{\infty}\lambda_{n}^{-1}\phi_{n}(x)\phi_{n}(x^{\prime}),\quad x,x^{\prime}\in\Omega.$ (3.3)
@@
Let $k$ be a symmetric, positive-definite kernel on $\Omega\times\Omega$, $|\Omega|<\infty$, satisfying $\int_{\Omega}k(x,x)dx<\infty$, and $S$ and $S^{-1}$ be the integral operators with kernels $k$ and $k^{-1}$, respectively. Let $u_{\theta}\in\mathrm{dom}(S^{-1})$ be parameterized by $\theta\in\mathbb{R}^{m}$. Then, a valid unbiased estimator for $\log p(\theta)$ as given by eq. (3.4) is
$\mathcal{S}_{N,M_{1},M_{2}}=-\frac{1}{2}\frac{|\Omega|^{2}}{NM_{1}M_{2}}\sum_{\alpha=1}^{N}\frac{1}{\lambda_{n_{\alpha}}p({n_{\alpha}})}\left\{\sum_{\beta=1}^{M_{1}}u_{\theta}(x_{\beta})\phi_{n_{\alpha}}(x_{\beta})\right\}\left\{\sum_{\gamma=1}^{M_{2}}u_{\theta}(x_{\gamma})\phi_{n_{\alpha}}(x_{\gamma})\right\},$ (3.5)
where $x_{\beta}\overset{\mathrm{i.i.d.}}{\sim}U(\Omega)$, $x_{\gamma}\overset{\mathrm{i.i.d.}}{\sim}U(\Omega)$, for $\beta=1,\dots,M_{1}$ and $\gamma=1,\dots,M_{2}$, and $n_{\alpha}$, $\alpha=1,\dots,N$, be i.i.d. draws from a discrete random variable with probability mass function $p(n)$, supported on $\mathbb{N}$. Here, $\lambda_{n_{\alpha}}$ and $\phi_{n_{\alpha}}$ represent the $n_{\alpha}$-th eigenvalue and corresponding eigenfunction of the operator $S$, respectively.
@@
$\log p(\theta)=-\frac{1}{2}\int_{\Omega}u(x)\int_{\Omega}k^{-1}(x,x^{\prime})u(x^{\prime})dx^{\prime}\>dx+\mathrm{const}.$
The assumptions on $k$ imply that $S$ is a trace-class operator, hence it is also a Hilbert-Schimdt operator. Therefore, the spectral theorem applies and the eigenvalues $\lambda_{n}$ and eigenfunctions $\phi_{n}$ exist. Also, each $\lambda_{n}$ is real and positive since $k$ is positive-definite.
@@
$k^{-1}(x,x^{\prime})=\sum_{n=1}^{\infty}\frac{1}{\lambda_{n}}\phi_{n}(x)\phi_{n}(x^{\prime}).$
@@
$\log p(\theta)$ $\overset{\Delta}{=}-\frac{1}{2}\int_{\Omega}u(x)\int_{\Omega}\sum_{n=1}^{\infty}\frac{1}{\lambda_{n}}\phi_{n}(x)\phi_{n}(x^{\prime})u(x^{\prime})dx^{\prime}\>dx$
$=-\frac{1}{2}\sum_{n=1}^{\infty}\frac{1}{\lambda_{n}}\int_{\Omega}u(x)\phi_{n}(x)dx\int_{\Omega}u(x^{\prime})\phi_{n}(x)^{\prime}dx^{\prime},$ (3.6)
@@
For the sampling averages, we employ importance sampling to write each piece of eq. (3) as an expectation. Firstly, if we let $X\sim U(\Omega)$, then we have
$\int_{\Omega}u_{\theta}(x)\phi_{n}(x)dt$ $=\int_{\Omega}\frac{|\Omega|}{|\Omega|}u_{\theta}(x)\phi_{n}(x)dt$
$=|\Omega|\mathbb{E}_{X\sim U(\Omega)}[u_{\theta}(X)\phi_{n}(X)]$
$\approx\frac{|\Omega|}{M_{1}}\sum_{\beta=1}^{M_{1}}u_{\theta}(x_{\beta})\phi_{n}(x_{\beta}),$
where $x_{\beta}$, $\beta=1,\dots,M_{1}$ are i.i.d. draws from $U(\Omega)$. This applies for both integrals.
@@
$\sum_{n=1}^{\infty}\frac{1}{\lambda_{n}}\int_{\Omega}u(x)\phi_{n}(x)dx^{\prime}\int_{\Omega}u(x^{\prime})\phi_{n}(x^{\prime})dx^{\prime}$ $=\sum_{n=1}^{\infty}\frac{p(n)}{\lambda_{n}p(n)}\langle u,\phi_{n}\rangle\langle u,\phi_{n}\rangle$

### v2

@@ Abstract
Quantifying the uncertainty in the output of a neural network is essential for deployment in scientific or engineering applications where decisions must be made under limited or noisy data. Bayesian neural networks (BNNs) provide a framework for this purpose by constructing a Bayesian posterior distribution over the network parameters. However, the prior, which is of key importance in any Bayesian setting, is rarely meaningful for BNNs. This is because the complexity of the input-to-output map of a BNN makes it difficult to understand how certain distributions enforce any interpretable constraint on the output space of the network. Gaussian processes (GPs), on the other hand, are often preferred in uncertainty quantification tasks due to their interpretability. The drawback is that GPs are limited to small datasets without advanced techniques, which often rely on the covariance kernel having a specific structure. To address these challenges, we introduce a new class of priors for BNNs, called Mercer priors, such that the resulting BNN has samples which approximate that of a specified GP. The method works by defining a prior directly over the network parameters from the Mercer representation of the covariance kernel, and does not rely on the network having a specific structure. In doing so, we can exploit the scalability of BNNs in a meaningful Bayesian way.
Keywords: Bayesian neural networks, Gaussian processes, surrogate modeling, inverse problems, reproducing kernel Hilbert spaces
@@ 1 Introduction
As neural networks begin to be employed in sensitive areas, such as complex engineering systems or healthcare [56, 45, 27, 21], the need to reliably assess their predictions becomes critically important. In these settings, it is oftentimes insufficient for a neural network to simply provide accurate point predictions, as the model must also quantify the uncertainty in the output to inform decisions. For neural networks, this task can be formulated under the Bayesian paradigm by treating the network parameters as random variables and constructing a posterior distribution conditioned on the data, which leads to a Bayesian neural network (BNN) [32, 36]. The complexity of the input-to-output map of a neural network makes it difficult to enforce any interpretable constraints on the output space of the BNN when assigning a specific form of the prior. As a result, the default choice is simply to place an independent and identically distributed Gaussian prior over the individual network parameters. Although this choice is certainly convenient, it is not always appropriate, and rarely translates to any meaningful constraints. In contrast, Gaussian processes (GPs) offer a greater degree of interpretability for uncertainty quantification tasks. This comes at a greater computational cost, however, as GPs do not scale well to large datasets without clever implementations that restrict the possible covariance structures.
Notably, a close relationship between BNNs and GPs is well-documented in the literature [29, 34, 38, 14]. For example, when the parameters of a BNN are drawn independently from a Gaussian distribution in the usual manner, the resulting distribution over functions converges to a GP in the infinite-width limit [36]. However, the form of the limiting GP depends entirely on the activation function of the network, as this determines the covariance kernel. This means that specifying a desired GP prior would require identifying a tailor-made activation function whose induced kernel matches that of the GP. While this has been shown to be possible for certain cases [61, 29, 66, 44], a general procedure remains out of reach. This observation motivates an alternative viewpoint: rather than fixing the activation function and sampling the parameters independently, is it instead possible to design the parameter distribution so that the BNN approximates a given GP?
@@
We first motivate the technique and contextualize the Mercer prior within the available methods for sampling BNNs in Sec. 2. The prior is introduced in Sec. 3, where we also derive a scalable sampling scheme and discuss possible covariance choices. After defining the Mercer prior, we do a deep dive into its use in sampling BNNs which resemble Brownian motion draws in Sec. 4. We compare the output statistics of the BNN to true Brownian motion, and provide numerical evidence that convergence is expected in the infinite-width limit. We then present three application-driven examples in Sec. 5: (i) hierarchical GP regression under heteroscedastic noise, (ii) time series prediction on data with a periodic structure, and (iii) a nonlinear elliptic PDE inverse problem with real-world data. Together, these results illustrate how Mercer priors provide a principled and scalable way to endow BNNs with an interpretable structure, enabling them to be deployed in uncertainty quantification tasks that are currently out of reach for standard GP models.
@@ 2 Background and setting
$u_{\theta}(t)=\sum_{j=1}^{N_{L}}w^{L}_{1,j}h\left(z_{j}^{L}(t)\right),\quad z_{i}^{s}(t)=\sum_{j=1}^{N_{s-1}}w_{i,j}^{s-1}h\left(z_{j}^{s-1}(t)\right)+b_{i}^{s-1},\quad s=2,\dots,L,$ (2.1)
where the input layer follows $z_{i}^{1}(t)=\sum_{j=1}^{d}w_{i,j}^{0}t_{j}+b_{i}^{0}$. We let $N_{s}=\dim(z^{s}(t))$ denote the number of neurons in layer $s$. The parameters of the BNN are the weights $w_{i,j}^{s}\in\mathbb{R}$ and the biases $b_{i}^{s}\in\mathbb{R}$, which are consolidated into the collection $\theta$. Eq. (2.1) defines a stochastic process $u:\Omega\times\Theta\to\mathbb{R}$, where $\Theta$ is a measurable parameter space on which a prior distribution on the BNN parameters $\theta$ is assigned. In the case $\dim(\Theta)<\infty$, i.e., for a finite-width network, we denote the prior over the BNN parameters as $p(\theta)$. We are also interested in studying infinite-width networks, where it is no longer rigorous to write $p(\theta)$. In this case, we denote the prior as some probability measure $\mu_{0}$ over the space $\Theta$. Either way, the mapping $\theta\mapsto u_{\theta}$ is a random field taking values in $L^{2}(\Omega)$.
@@
which holds even when $\dim(\Theta)=\infty$ [51]. In the case of a finite-width network, this reduces to the well-known formula $p(\theta|y)\propto p(y|\theta)p(\theta)$, with the likelihood being $p(y|\theta)\propto\exp\left(-\tfrac{1}{2}\|\Gamma^{-1}(y-R(u_{\theta}))\|^{2}\right)$. Note that this form of the posterior arises under Gaussian measurement noise. In the finite-dimensional case, as with training a BNN, non-Gaussian noise is easily handled by a straightforward application of Bayes’s theorem.
@@ 2.1 Wide neural networks and Gaussian processes
$k(t,t^{\prime})=\sigma_{w}^{2}\mathbb{E}[h(t^{\intercal}w_{1}^{0}+b_{1}^{0})h(t^{\prime\intercal}w_{1}^{0}+b_{1}^{0})]+\sigma_{b}^{2}.$
@@
$u(\cdot)=\sum_{n=1}^{\infty}\sqrt{\lambda_{n}}\xi_{n}\phi_{n}(\cdot),$
where $\lambda_{n}$ and $\phi_{n}$ are the eigenvalues and eigenfunctions of the kernel integral operator $Su=\int k(\cdot,t)u(t)dt$, and $\xi_{n}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,1)$. The expansion is truncated to a finite number of terms $M$ in practice. When using the KLE, the computational complexity is transferred from inverting a large matrix to constructing the basis functions. However, since the eigenfunctions $\phi_{n}$ are global and fixed, performance may suffer when the data is non-uniformly spaced or sparse, unless $M$ is large. Also, the KLE suffers from the curse of dimensionality, as the number of eigenfunctions needed to capture most of the variance grows exponentially with the dimension. A comprehensive review of these ideas can be found in [30].
@@ 2.2 Related methods
Looking to gain both the expressivity of GPs and scalability of BNNs, several methods have been proposed that exploit the relationship between the two. This can be done using the correspondence between a BNN and the associated NNGP, so that the problem of prior selection for a BNN is transferred to the easier problem of prior selection for GPs. One class of methods selects the prior in order to minimize some objective between the BNN and the target GP [18, 20, 52]. The drawback is that the user is constrained to a particular algorithm for posterior inference (like variational inference). These methods also typically need to see some of the training data upfront in order to construct the prior, e.g. [35]. Other methods rely on changing the architecture of the BNN to ensure convergence to the GP in an appropriate sense [41, 2, 44]. This of course works on a case-by-case basis.
@@
The ridgelet prior and the Mercer prior we present in this work are attempting to solve the same problem. Namely that we want to sample the parameters of a BNN such that the resulting function resembles draws from a predetermined GP. The difference then comes down to how the prior over the network parameters is characterized. In this sense, one can view the two as different finite-dimensional approximations of the same GP. However, computing the Mercer prior scales linearly in the number of network parameters, which allows for the potential use of larger networks. While the Mercer prior does contain two integrals which need to be computed, we avoid the curse of dimensionality by deriving a sampling scheme which uses an unbiased estimate of the prior. This is built with Monte Carlo estimates, which in principle guarantees convergence even for a batch size of $1$ (this is equivalent to setting $D=1$ in the ridgelet prior).
@@ 3 The Mercer prior
Start by defining a GP on $\Omega\subset\mathbb{R}^{d}$ that we wish to emulate $u\sim\mathcal{GP}(0,k)$, where $k:\Omega\times\Omega\to\mathbb{R}$ is the covariance kernel. Note that we are choosing to work with a centered GP to simplify some of the details. A non-zero mean function could be used instead, provided that it satisfies certain regularity constraints. The resulting formulae would then need to be adjusted. Alternatively, one could shift the space so that the GP becomes centered. Further, for the covariance kernel, we enforce the restriction $\int_{\Omega}k(t,t)dt<\infty$. This is so that we may use the following theorem, which relates the GP to an induced Gaussian measure on $L^{2}(\Omega)$. A positive-definite kernel statisfying these properties is sometimes called a Mercer kernel.
@@
$\int_{\Omega}m^{2}(t)dt<\infty,\quad\int_{\Omega}k(t,t)dt<\infty.$
In this case, $u$ induces the Gaussian measure $\mathcal{N}(m,S)$ on $L^{2}(\Omega)$ with the covariance operator being $(Sv)(\cdot)=\int_{\Omega}k(\cdot,t)v(t)dt$, for $v\in L^{2}(\Omega)$.
@@
We then numerically characterize eq. (3.2) using traditional Monte Carlo sampling schemes in order sample BNNs. Each sample of $\theta$ provides a new neural network, and we will see that the ensemble statistics approximately follow that of the original Gaussian measure $\mu$. This offers a significant reduction in computational complexity when the BNN sampled from eq. (3.2) replaces a GP for the same task. As mentioned previously, the complexity of a GP scales with the square of the evaluation mesh. On the other hand, the BNN prior of eq. (3.2) grows with $\theta$. After sampling, the BNN can be evaluated on a domain with a near-arbitrarily fine mesh, which allows for easy super-resolution.
@@
$(S^{-1}v)(\cdot)=\int_{\Omega}k^{-1}(\cdot,t)v(t)dt,\quad v\in L^{2}(\Omega),$
such that $S^{-1}(Sv)=v$.
We work our way around the need to evaluate the integrals by deriving a sampling scheme which uses stochastic gradient Langevin dynamics (SGLD) [58]. The beauty of SGLD is that samples from $p(\theta)$ can be generated using only unbiased estimates of $\nabla_{\theta}\log p(\theta)=-1/2\nabla_{\theta}\langle u_{\theta},S^{-1}u_{\theta}\rangle$. This allows us to approximate the various integrals required using sampling averages and SGLD guarantees convergence to the correct probability distribution [54].
@@
$(Su)(\cdot)=\sum_{n=1}^{\infty}\lambda_{n}\langle u,\phi_{n}\rangle\phi_{n}(\cdot),\quad u\in L^{2}(\Omega),$
@@
Let $k:\Omega\times\Omega\to\mathbb{R}$ be a continuous, positive-definite kernel, and $S$ be the operator on $L^{2}(\Omega)$ defined by $u\mapsto\int_{\Omega}k(\cdot,t)u(t)dt.$ Denote the eigenvalues and eigenfunctions of $S$ as $\{\lambda_{n}\}_{n\in\mathbb{N}}$ and $\{\phi_{n}\}_{n\in\mathbb{N}}$, respectively. Then $k$ can be expressed as
$k(s,t)=\sum_{n=1}^{\infty}\lambda_{n}\phi_{n}(s)\phi_{n}(t),\quad s,t\in\Omega,$
@@
$k^{-1}(s,t)=\sum_{n=1}^{\infty}\lambda_{n}^{-1}\phi_{n}(s)\phi_{n}(t),\quad s,t\in\Omega.$ (3.3)
@@
Let $k$ be a symmetric, positive-definite kernel on $\Omega\times\Omega$, $|\Omega|<\infty$, satisfying $\int_{\Omega}k(t,t)dt<\infty$, and $S$ and $S^{-1}$ be the integral operators with kernels $k$ and $k^{-1}$, respectively. Let $u_{\theta}\in\mathrm{dom}(S^{-1})$ be parameterized by $\theta\in\mathbb{R}^{m}$. Then, a valid unbiased estimator for $\log p(\theta)$ as given by eq. (3.4) is
$\mathcal{S}_{N,M_{1},M_{2}}=-\frac{1}{2}\frac{|\Omega|^{2}}{NM_{1}M_{2}}\sum_{\alpha=1}^{N}\frac{1}{\lambda_{n_{\alpha}}p({n_{\alpha}})}\left\{\sum_{\beta=1}^{M_{1}}u_{\theta}(t_{\beta})\phi_{n_{\alpha}}(t_{\beta})\right\}\left\{\sum_{\gamma=1}^{M_{2}}u_{\theta}(t_{\gamma})\phi_{n_{\alpha}}(t_{\gamma})\right\},$ (3.5)
where $t_{\beta}\overset{\mathrm{i.i.d.}}{\sim}U(\Omega)$, $t_{\gamma}\overset{\mathrm{i.i.d.}}{\sim}U(\Omega)$, for $\beta=1,\dots,M_{1}$ and $\gamma=1,\dots,M_{2}$, and $n_{\alpha}$, $\alpha=1,\dots,N$, be i.i.d. draws from a discrete random variable with probability mass function $p(n)$, supported on $\mathbb{N}$. Here, $\lambda_{n_{\alpha}}$ and $\phi_{n_{\alpha}}$ represent the $n_{\alpha}$-th eigenvalue and corresponding eigenfunction of the operator $S$, respectively.
@@
$\log p(\theta)=-\frac{1}{2}\int_{\Omega}u(s)\int_{\Omega}k^{-1}(s,t)u(t)dt\>ds+\mathrm{const}.$
The assumptions on $k$ imply that $S$ is a trace-class operator, hence it is also a Hilbert-Schmidt operator. Therefore, the spectral theorem applies and the eigenvalues $\lambda_{n}$ and eigenfunctions $\phi_{n}$ exist. Also, each $\lambda_{n}$ is real and positive since $k$ is positive-definite.
@@
$k^{-1}(s,t)=\sum_{n=1}^{\infty}\frac{1}{\lambda_{n}}\phi_{n}(s)\phi_{n}(t).$
@@
$\log p(\theta)$ $\overset{\Delta}{=}-\frac{1}{2}\int_{\Omega}u(s)\int_{\Omega}\sum_{n=1}^{\infty}\frac{1}{\lambda_{n}}\phi_{n}(s)\phi_{n}(t)u(t)dt\>ds$
$=-\frac{1}{2}\sum_{n=1}^{\infty}\frac{1}{\lambda_{n}}\int_{\Omega}u(s)\phi_{n}(s)ds\int_{\Omega}u(t)\phi_{n}(t)dt,$ (3.6)
@@
For the sampling averages, we employ importance sampling to write each piece of eq. (3) as an expectation. Firstly, if we let $T\sim U(\Omega)$, then we have
$\int_{\Omega}u_{\theta}(t)\phi_{n}(t)dx$ $=\int_{\Omega}\frac{|\Omega|}{|\Omega|}u_{\theta}(t)\phi_{n}(t)dt$
$=|\Omega|\mathbb{E}_{T\sim U(\Omega)}[u_{\theta}(T)\phi_{n}(T)]$
$\approx\frac{|\Omega|}{M_{1}}\sum_{\beta=1}^{M_{1}}u_{\theta}(t_{\beta})\phi_{n}(t_{\beta}),$
where $t_{\beta}$, $\beta=1,\dots,M_{1}$ are i.i.d. draws from $U(\Omega)$. This applies for both integrals.
@@
$\sum_{n=1}^{\infty}\frac{1}{\lambda_{n}}\int_{\Omega}u(s)\phi_{n}(s)ds\int_{\Omega}u(t)\phi_{n}(t)dt$ $=\sum_{n=1}^{\infty}\frac{p(n)}{\lambda_{n}p(n)}\langle u,\phi_{n}\rangle\langle u,\phi_{n}\rangle$